OpenStudy (anonymous):

Express the sum or difference as a product. sin 4x - sin 6x A. 2 cos 4x cos 5x B. 2 sin 5x cos x C. -2 sin x cos 5x D. -2 sin x

OpenStudy (anonymous):

@jim_thompson5910 @radar @sami-21

OpenStudy (anonymous):

@cwrw238

jimthompson5910 (jim_thompson5910):

Use the identity \[\Large \sin(A)-\sin(B) = 2\cos\left(\frac{A+B}{2}\right)\sin\left(\frac{A-B}{2}\right)\]

OpenStudy (cwrw238):

this is a sum-to-product identity you can find them here http://www.sosmath.com/trig/Trig5/trig5/trig5.html oh - jim s got it

OpenStudy (anonymous):

I got 2 cos(5x) sin(-x) that not a choice @jim_thompson5910

OpenStudy (cwrw238):

we used to try remembering these by something like 'twice cos semi sum sine semi difference' - never found that very helpful

jimthompson5910 (jim_thompson5910):

sine is an odd function, which means sin(-x) = -sin(x)

OpenStudy (anonymous):

ohh so C?

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

thanks!

jimthompson5910 (jim_thompson5910):

np

OpenStudy (radar):

@cwrw238, thanks for the link.

OpenStudy (cwrw238):

thats ok